OpenStudy (anonymous):

Medals shall be rewarded. Thanks for the help (Trigonometry question) A Ferris wheel with radius 31m makes 2 rotations in one minute. a) Find the angular velocity in radians per second. b) How far has a rider traveled if the ride is 10 minutes long? (I basically converted the 2 revolutions per minute in 0.21 rad/sec as the final answer for part (a), which is correct, but I am having trouble solving part (b)).

OpenStudy (mertsj):

It makes 1 rotation in 30 seconds. That is 2pi radians in 30 seconds. So it goes 4 pi radians in 60 seconds. So a is: \[\frac{4\pi}{60}=\frac{\pi}{15} radians per second\]

OpenStudy (mertsj):

b) If the ferris wheel has a radius of 31 m, you can find its circumference. What is it?

OpenStudy (anonymous):

According to the book I have, the distance is 974m

OpenStudy (mertsj):

\[C=2\pi r=2(3.14)(31)=194.68 m\] in 1 revolution. So in 2 revolutions (1 minute) it goes 389.36 m. And in 10 minutes, it goes 10(389.36) of 3893.6 m

OpenStudy (anonymous):

That is the very answer I got as well. Maybe the book is wrong?

OpenStudy (mertsj):

Either that or you have posted part b incorrectly.

OpenStudy (anonymous):

Okay. Thank you very much.

OpenStudy (mertsj):

Oh...you could double check the units. Are you sure it said 974 meters?

OpenStudy (anonymous):

Yup. 974m.

OpenStudy (mertsj):

Then it is incorrect.

OpenStudy (anonymous):

Okay. Thanks again.

OpenStudy (mertsj):

http://msmklein.wikispaces.com/file/view/2.Angular+velocity.pdf Scroll down to page 3 and look at problem number 2. The solution is on page 4

OpenStudy (anonymous):

Okay, the copy I received definitely has errors in it.

OpenStudy (mertsj):

yep.